### v1

# Delayed Acceptance Markov Chain Monte Carlo for Robust Bayesian Analysis

Abstract
This study introduces a computationally efficient algorithm, delayed acceptance Markov chain Monte Carlo (DA–MCMC), designed to improve posterior simulation in quasi-Bayesian inference. Quasi-Bayesian methods, which do not require fully specifying a probabilistic model, are often computationally expensive owing to the need to evaluate the inverse and determinant of large covariance matrices. DA–MCMC addresses this challenge by employing a two-stage process: In the first stage, proposals are screened using an approximate posterior, whereas a final acceptance or rejection decision is made in the second stage based on the exact target posterior. This reduces the need for costly matrix computations, thereby improving efficiency without sacrificing accuracy. We demonstrate the effectiveness of DA–MCMC through applications to both synthetic and real data. The results demonstrate that, although DA–MCMC slightly reduces the effective sample size per iteration compared with the standard MCMC, it achieves substantial improvement in terms of effective sample size per second, approximately doubling the efficiency. This makes DA–MCMC particularly useful for cases where posterior simulation is computationally intensive. Thus, the DA–MCMC algorithm offers a significant advancement in computational efficiency for quasi-Bayesian inference, making it a valuable tool for robust Bayesian analysis.
Keywords: delayed acceptance Markov chain Monte Carlo; robust Bayesian analysis; generalized posterior; Gibbs posterior
1 Introduction
In a standard Bayesian inference, a probabilistic model is used for the data-generating process and the corresponding likelihood function, $p\left(\mathcal{D}|\boldsymbol{\theta}\right)$, where $\mathcal{D}$ denotes a set of observations with a sample size of $N$, $\boldsymbol{\theta}\in\mathbb{R}^{K}$ denotes a $K$-dimensional vector of unknown parameters, and $p\left(\boldsymbol{\theta}\right)$ denotes a prior. A posterior density is proportional to the product of the likelihood and prior, $p\left(\boldsymbol{\theta}|\mathcal{D}\right)\propto p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)$. Because a posterior is generally intractable, we simulate random draws from the density using numerical techniques, such as Markov chain Monte Carlo (MCMC) algorithms.
A well-known shortcoming of standard Bayesian analysis is that it needs to fully specify a probabilistic model. Alternative quasi-Bayesian approaches have been proposed to improve the robustness toward model misspecification, which have different constructions of quasi-likelihood that rely on loss function and moment conditions instead of a probabilistic model.
This study considers two types of quasi-Bayesian approaches. In the first approach, a quasi-likelihood is based on an empirical risk function $r\left(\boldsymbol{\theta};\mathcal{D}\right)$ (e.g., [19, 20, 9, 4, 13]). Hereinafter, we ignore the dependency on $\mathcal{D}$ and write $r\left(\boldsymbol{\theta}\right)=r\left(\boldsymbol{\theta};\mathcal{D}\right)$. For instance, consider a linear regression with response variable $y_{i}$, covariates $\boldsymbol{x}_{i}$, and coefficients $\boldsymbol{\theta}$. The unknown parameter $\boldsymbol{\theta}$ can be estimated using an empirical risk function with quadratic losses, $r\left(\boldsymbol{\theta}\right)=N^{-1}\sum_{i=1}^{N}\left(y_{i}-\boldsymbol{\theta}^{\top}\boldsymbol{x}_{i}\right)^{2}$ where $N$ denotes a sample size. Here, a probabilistic assumption does not have to be imposed on the error term such as the normality; thus, $\boldsymbol{\theta}$ can be estimated with more robustness. Given an empirical risk function, a Gibbs posterior is defined as follows:
$\pi_{\omega}\left(\boldsymbol{\theta}\right)\propto\left[\exp\left(-Nr\left(\boldsymbol{\theta}\right)\right)\right]^{\omega}p\left(\boldsymbol{\theta}\right).$ (1)
In this approach, a scaling parameter (also known as a learning rate parameter) $\omega\left(>0\right)$ must be selected to calibrate the spread of the posterior; however, it is not a theoretically or computationally trivial problem. [^1] See [16] for a review of the related literature. Recently, Frazier et al. [7] proposed a new approach that specifies a quasi-posterior as follows:
$\pi\left(\boldsymbol{\theta}\right)\propto\left|\boldsymbol{W}\left(\boldsymbol{\theta}\right)\right|^{-\frac{1}{2}}\exp\left\{-\frac{N}{2}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)\right\}p\left(\boldsymbol{\theta}\right),$ (2)
where $\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)=\nabla r\left(\boldsymbol{\theta}\right)$, $\nabla$ denotes the gradient operator, and $\boldsymbol{W}\left(\boldsymbol{\theta}\right)$ denotes an estimate of the covariance of $\sqrt{N}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)$. As argued in [7], this approach can provide proper uncertainty quantification without selecting $\omega$ under some regularity conditions.
The second approach specifies a quasi-likelihood based on a generalized method of moments (GMM) criterion (e.g., [10, 5, 17]). In a GMM, a statistical problem is formulated under a set of moment conditions, $E\left[\boldsymbol{m}_{i}\left(\boldsymbol{\theta};\mathcal{D}\right)\right]=\boldsymbol{0}$, where $\boldsymbol{m}_{i}\left(\cdot\right)$ denotes a vector-valued function termed a moment function. We shortly denote $\boldsymbol{m}_{i}\left(\boldsymbol{\theta}\right)=\boldsymbol{m}_{i}\left(\boldsymbol{\theta};\mathcal{D}\right)$. The unknown parameters $\boldsymbol{\theta}$ are estimated by minimizing a GMM criterion $\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)$, with $\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)=N^{-1}\sum_{i=1}^{N}\boldsymbol{m}_{i}\left(\boldsymbol{\theta}\right)$. Then, a quasi-posterior is given by
$\pi_{\omega}\left(\boldsymbol{\theta}\right)\propto\left[\left|\boldsymbol{W}\left(\boldsymbol{\theta}\right)\right|^{-\frac{1}{2}}\exp\left\{-\frac{N}{2}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)\right\}\right]^{\omega}p\left(\boldsymbol{\theta}\right).$ (3)
This approach can be regarded as a special case of (2) when the moment conditions are derived under the first-order condition for an empirical risk function. Although the first approach can only handle exactly identified models, the second can also handle over-identified models. For notational convenience, we subsequently focus on (2).
The biggest challenge in implementing Bayesian inference with (2) is the computational load. The evaluation of $\pi\left(\boldsymbol{\theta}\right)$ is computationally expensive because of the inverse matrix $\boldsymbol{W}\left(\boldsymbol{\theta}\right)^{-1}$ as well as the matrix determinant $\left|\boldsymbol{W}\left(\boldsymbol{\theta}\right)\right|$. To the best of our knowledge, no studies have addressed this issue. Yin et al. [18] proposed a posterior simulator for inference with (3); however, their proposal aimed to address the numerical stability and can be computationally expensive.
In this study, we propose a new algorithm based on a delayed acceptance MCMC (DA–MCMC) [6], which can be beneficial for Monte Carlo simulation when the target kernel is expensive to compute. DA–MCMC can be more computationally efficient than the standard MCMC because it can shorten the computation of $\boldsymbol{W}\left(\boldsymbol{\theta}\right)^{-1}$. A single cycle of DA–MCMC has two stages. In the first stage, after being generated for a new state, the proposal is promoted to the second stage with probability based on an approximate target density. In the second stage, the promoted proposal is accepted with probability based on the exact target density. Through applications to synthetic and real data problems, we demonstrate that the proposed algorithm is approximately two times more efficient than the standard MCMC in terms of the effective sample size per second.
The remainder of this paper is organized as follows: Section 2 introduces DA–MCMC, Section 3 compares the proposed algorithm with existing algorithm via synthetic and real data examples, and Section 4 concludes the paper.
2 Delayed acceptance MCMC
To expedite the posterior simulation with (2), we propose using DA–MCMC [6]. A single cycle of the algorithm comprises two stages: (1) screening a proposal with an approximate posterior and (2) determining whether to accept or reject with the exact posterior. The second stage occurs only when a proposal passes the first stage.
Given a current state $\boldsymbol{\theta}^{\left(t\right)}$, a proposal $\boldsymbol{\theta}^{\prime}$ is generated from a proposal kernel, $\boldsymbol{\theta}^{\prime}\sim q_{1}\left(\cdot|\boldsymbol{\theta}^{\left(t\right)}\right)$. The first-stage acceptance probability is based on an approximation to the exact target kernel $\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)$, which is evaluated with less computational cost than $\pi\left(\boldsymbol{\theta}\right)$. An approximation to the target kernel is constructed by replacing $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)$ with the current value $\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)$:
$\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)=\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)\right|^{-\frac{1}{2}}\exp\left\{-\frac{N}{2}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}^{\prime}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}^{\prime}\right)\right\}p\left(\boldsymbol{\theta}^{\prime}\right).$ (4)
With this approximation, we do not need to compute the most computationally demanding quantities, that is, the inverse and determinant of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)$. The first-stage acceptance probability is computed as follows:
$\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{1}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)}{q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi^{*}\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$ (5)
With probability $\alpha_{1}\left(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}^{\left(t\right)}\right)$, $\boldsymbol{\theta}^{\prime}$ is “promoted” as a proposal for the second stage; otherwise, $\boldsymbol{\theta}^{\prime}$ is rejected, thereby maintaining the current state $\boldsymbol{\theta}^{\left(t+1\right)}=\boldsymbol{\theta}^{\left(t\right)}$. In the second stage, the acceptance probability is evaluated based on the exact target kernel $\pi\left(\boldsymbol{\theta}^{\prime}\right)$ as follows:
$\alpha_{2}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{2}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi\left(\boldsymbol{\theta}^{\prime}\right)}{q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$ (6)
where $q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)$ denotes the actual proposal density given by:
$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right).$ (7)
As argued in [6], compared with the standard MCMC, a single cycle of DA–MCMC can be slightly less efficient in terms of effective sample size per iteration. However, DA–MCMC can be more efficient in terms of the effective sample size per second because it does not have to perform the unnecessary computation of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)^{-1}$ and $\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)\right|$. As shown in the next section, DA–MCMC is effective in our context.
This study uses an adaptive random walk Metropolis–Hastings kernel, $q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\mathrm{N}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)},\;\varepsilon\boldsymbol{\Sigma}\right)$, where $\varepsilon>0$ denotes a step size, $\boldsymbol{\Sigma}$ denotes the $K$-dimensional symmetric positive definite matrix, and $\mathrm{N}\left(\boldsymbol{x}|\boldsymbol{a},\boldsymbol{B}\right)$ denotes a probability density function of a multivariate normal distribution with mean $\boldsymbol{a}$ and covariance matrix $\boldsymbol{B}$ evaluated at $\boldsymbol{x}$. The parameters of the proposal kernel are adaptively selected on the fly of an MCMC algorithm [8]. At the $t$-th cycle, $\varepsilon$ is updated as $\log\varepsilon\leftarrow\log\varepsilon+t^{-\varsigma}\left(\bar{\alpha}^{\left(t\right)}-\alpha^{\dagger}\right)$, where $\bar{\alpha}^{\left(t\right)}$ denotes the average acceptance probability up to the $t$-th cycle, $\alpha^{\dagger}\in\left(0,1\right)$ denotes the target acceptance probability, and $\varsigma\in\left(0.5,1\right)$ denotes a tuning parameter. In this study, we fix $\varsigma=0.51$. $\boldsymbol{\Sigma}$ is updated based on the empirical covariance matrix of $\boldsymbol{\theta}^{\left(1\right)},\boldsymbol{\theta}^{\left(2\right)},...,\boldsymbol{\theta}^{\left(t\right)}$. The adaptation of $\varepsilon$ and $\boldsymbol{\Sigma}$ is performed only in the warmup phase and not thereafter.
3 Application
3.1 Synthetic data: Heteroskedastic linear regression
We considered a heteroskedastic linear regression with synthetic data. The data-generating process was adopted from [7]. A response variable $y_{i}$ was generated from a normal distribution, $y_{i}\sim\mathcal{N}\left(\boldsymbol{\theta}^{\top}\boldsymbol{x}_{i},\sigma_{i}^{2}\right)$, where a vector of covariates $\boldsymbol{x}_{i}=\left(x_{1,i},x_{2,i},...,x_{K,i}\right)^{\top}$ with $x_{1,i}=1$, $\boldsymbol{\theta}\in\mathbb{R}^{K}$ denotes a vector of coefficients, and the variance is specified as $\sigma_{i}^{2}=\left(1+\left|x_{2,i}\right|^{2}+\left|x_{3,i}\right|^{2}\right)/3$. The true values were set to $\boldsymbol{\theta}=\left(1,1,1,0,...,0\right)^{\top}$. The covariates except $x_{1,i}$ were simulated from the independent standard normal distribution. The unknown parameters were estimated via an empirical risk function with quadratic losses, $r\left(\boldsymbol{\theta}\right)=N^{-1}\sum_{i=1}^{N}\left(y_{i}-\boldsymbol{\theta}^{\top}\boldsymbol{x}_{i}\right)^{2}$. The prior of $\boldsymbol{\theta}$ was specified as $\boldsymbol{\theta}\sim\mathcal{N}\left(\boldsymbol{0}_{K},100^{2}\boldsymbol{I}_{K}\right)$.
We considered four settings with combinations of $N\in\left\{100,1000\right\}$ and $K\in\left\{5,20\right\}$. In each posterior simulation, a total of 20,000 draws were generated and the last 10,000 draws were used for evaluation. The numerical performance was evaluated based on multivariate effective sample size (multiESS) [14].  [^2] The programs were executed on MATLAB (R2024b) on an Ubuntu desktop (22.04.5 LTS) running on an AMD Ryzen Threadripper 3990X 2.9 GHz 64-core processor.
The choice of the target acceptance rate $\alpha^{\dagger}$ is of practical importance. However, no specific theoretical guidance on this matter exists, as the theoretical properties of DA–MCMC algorithms have not been investigated thoroughly. To address this gap, we empirically examine the choice of $\alpha^{\dagger}$ using a simulation study. Figure 1 displays the medians of multiESS/iter for different target acceptance rates $\alpha^{\dagger}\in\left\{0.02,0.04,...,0.4\right\}$ based on 1,000 independent runs for each setting. Except for the case with $N=100$ and $K=20$, multiESS/iter is maximized approximately 0.2 to 0.3. Thus, adapting the DA–MCMC with a random walk Metropolis–Hastings steps to achieve an acceptance rate of 25% ($\alpha^{\dagger}=0.25$) can be recommended as a default option. However, selecting $\alpha^{\dagger}=0.25$ can be suboptimal in certain cases, such as when $N$ is insufficiently large relative to $K$. As an experiment, we set $\alpha^{\dagger}=0.05$ and ran DA–MCMC on the synthetic data with $N=100$ and $K=20$ to obtain an increased multiESS/s of 1,515 (compared with 993 having $\alpha^{\dagger}=0.25$). Further investigation into the optimal selection of $\alpha^{\dagger}$ will be the focus of future research.
Figure 1: MultiESS/iter for different target acceptance rate
Tables 1 and 2 list the medians of the multiESS per iteration (multiESS/iter) and multiESS per second (multiESS/s) for 1,000 runs, respectively. As expected, the multiESS/iter for DA–MCMC tended to be slightly smaller than that for the standard MCMC; however, the difference was small. In contrast, the multiESS/s for DA–MCMC was approximately 1.5 to 2 times larger than that for the standard MCMC. As shown in Table 1, the absolute level of efficiency per iteration for this type of algorithm is low, requiring the generation of long chains to converge. Therefore, improving efficiency per unit time is an important contribution.
Table 1: Multivariate effective sample size per iteration
Model $N$ $K$ multiESS/iter
MCMC DA–MCMC
Linear Reg. 100 5 0.053 0.045
20 0.014 0.014
1000 5 0.062 0.062
20 0.023 0.022
IV Reg. 64 6 0.020 0.021
Table 2: Multivariate effective sample size per second
Model $N$ $K$ multiESS/s
MCMC DA–MCMC
Linear Reg. 100 5 4,064 6,449
20 467 993
1000 5 2,494 4,834
20 204 515
IV Reg. 64 6 160 293
To further investigate the behavior of DA–MCMC, Figure 2 presents the relative frequency of the second-stage acceptance probability for linear regression using synthetic data. Table 3 summarizes the results of the 25th, 50th, and 75th percentiles of these probabilities. Except for the case with $N=100$ and $K=20$, most of the observed second-stage acceptance probabilities were close to 1. This indicates that once a proposal passes the first-stage screening, it is almost certain that it will be considered acceptable in the second stage. However, in the most challenging task, where $N=100$ and $K=20$ (where $K$ is large relative to $N$), the second-stage acceptance probabilities were widely dispersed and ranged from 0 to 1. The median second-stage acceptance probability was 0.250, suggesting that most of the promoted proposals were rejected. This implies that the first-stage posterior was not a good approximation, because the contributions of $\boldsymbol{W}\left(\boldsymbol{\theta}\right)^{-1}$ and $\left|\boldsymbol{W}\left(\boldsymbol{\theta}\right)\right|$ to the (true) posterior were likely significant.
Figure 2: Relative frequency of second-stage acceptance probability
Table 3: Second-stage acceptance probability
$N$ $K$ Percentile
$P_{25}$ $P_{50}$ $P_{75}$
100 5 0.760 0.958 1.000
20 0.062 0.250 0.798
1000 5 0.978 0.999 1.000
20 0.907 0.992 1.000
3.2 Real data: Instrumental variable regression
We estimated an instrumental variable regression using a real dataset. The dataset is obtained from Acemoglu et al. [2, 3]. [^3] The dataset is downloaded from the OPENICPSR website. https://www.openicpsr.org/openicpsr/project/112564/version/V1/view They investigated the effect of the expropriation risk ($x_{i}$) on the gross domestic product per capita (in logs) ($y_{i}$). They used the European settler mortality (in logs) as an instrument ($z_{i}$) to address endogeneity. The vector of controls $\boldsymbol{w}_{i}$ includes a constant, latitude, and continents dummies, namely, dummies for Africa, Asia, and the former British colonies (Australia, Canada, New Zealand, and the United States).
The model is expressed as follows:
$x_{i}$ $=$ $g\left(z_{i},\boldsymbol{w}_{i}\right)+v_{i},$
$y_{i}$ $=$ $\mu+\beta x_{i}+\boldsymbol{\gamma}^{\top}\boldsymbol{w}_{i}+u_{i},$
where $v_{i}$, and $u_{i}$ denote error terms. Here, $g\left(\cdot\right)$ denotes an unknown function. The unknown parameters $\boldsymbol{\theta}=\left(\mu,\beta,\boldsymbol{\gamma}^{\top}\right)^{\top}$ are inferred under the following moment conditions:
$\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)=\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}-\left(\mu+\beta x_{i}+\boldsymbol{\gamma}^{\top}\boldsymbol{w}_{i}\right)\right)\left(\begin{array}[]{c}z_{i}\\ \boldsymbol{w}_{i}\end{array}\right).$
For $\boldsymbol{\theta}$, we use a normal prior, $\boldsymbol{\theta}\sim\mathcal{N}\left(\boldsymbol{0}_{K},100^{2}\boldsymbol{I}_{K}\right)$. This approach has notable advantages over standard Bayesian approaches to instrumental variable regression, such as [11, 15, 12]. First, it does not require the estimation of $g\left(\cdot\right)$. Second, it does not specify the distributions of $v_{i}$ and $u_{i}$. The sample size is $N=64$ and the number of unknown parameters is $K=6$. We discarded the first 100,000 draws and used the last 1,000,000 draws for analysis.
The comparison of the numerical performance based on 1,000 runs is summarized in the last rows of Tables 1 and 2. As in the application to synthetic data above, although the multiESS/iter for DA–MCMC was virtually indistinct from that for the standard MCMC, the multiESS/s for DA–MCMC was much larger than that for the standard MCMC. Figure 3 compares the simulated posterior distributions of $\beta$ using the two algorithms. The difference between the two was virtually negligible. The result shows that these two algorithms can generate random numbers from the same distribution, despite their significantly different computational efficiencies.
Figure 3: Posterior distribution of $\beta$
4 Conclusion
This study introduced a computationally efficient algorithm for Bayesian inference, specifically tailored for quasi-Bayesian approaches using DA–MCMC. By leveraging the two-stage screening and promotion process of DA–MCMC, we demonstrated that it significantly reduced the computational burden of evaluating the quasi-posterior, particularly when considering high-dimensional models and costly matrix computations. Through synthetic and real data applications, we demonstrated that although the DA–MCMC algorithm slightly decreased effective sample size per iteration compared with the standard MCMC, it offered a substantial improvement in terms of computational efficiency, approximately doubling the effective sample size per second. This efficiency gain holds promise for broader applications in complex statistical models where posterior simulation can otherwise be prohibitively slow. Future work will explore optimizing the algorithm further for large-scale problems or its application in other quasi-Bayesian frameworks.
Acknowledgements
This work was supported by JSPS KAKENHI Grant Number 20K22096.

### v2

@@
 Thus, the DA–MCMC algorithm offers a significant advancement in computational efficiency for quasi-Bayesian inference, making it a valuable tool for robust Bayesian analysis.
-Keywords: delayed acceptance Markov chain Monte Carlo; robust Bayesian analysis; generalized posterior; Gibbs posterior
+…
+keywords:
+delayed acceptance Markov chain Monte Carlo; robust Bayesian analysis; generalized posterior; Gibbs posterior
 1 Introduction
@@
 2 Delayed acceptance MCMC
-To expedite the posterior simulation with (2), we propose using DA–MCMC [6].
-A single cycle of the algorithm comprises two stages: (1) screening a proposal with an approximate posterior and (2) determining whether to accept or reject with the exact posterior.
-The second stage occurs only when a proposal passes the first stage.
-Given a current state $\boldsymbol{\theta}^{\left(t\right)}$, a proposal $\boldsymbol{\theta}^{\prime}$ is generated from a proposal kernel, $\boldsymbol{\theta}^{\prime}\sim q_{1}\left(\cdot|\boldsymbol{\theta}^{\left(t\right)}\right)$.
-The first-stage acceptance probability is based on an approximation to the exact target kernel $\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)$, which is evaluated with less computational cost than $\pi\left(\boldsymbol{\theta}\right)$.
-An approximation to the target kernel is constructed by replacing $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)$ with the current value $\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)$:
-$\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)=\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)\right|^{-\frac{1}{2}}\exp\left\{-\frac{N}{2}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}^{\prime}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}^{\prime}\right)\right\}p\left(\boldsymbol{\theta}^{\prime}\right).$ (4)
-With this approximation, we do not need to compute the most computationally demanding quantities, that is, the inverse and determinant of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)$.
-The first-stage acceptance probability is computed as follows:
-$\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{1}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)}{q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi^{*}\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$ (5)
-With probability $\alpha_{1}\left(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}^{\left(t\right)}\right)$, $\boldsymbol{\theta}^{\prime}$ is “promoted” as a proposal for the second stage; otherwise, $\boldsymbol{\theta}^{\prime}$ is rejected, thereby maintaining the current state $\boldsymbol{\theta}^{\left(t+1\right)}=\boldsymbol{\theta}^{\left(t\right)}$.
-In the second stage, the acceptance probability is evaluated based on the exact target kernel $\pi\left(\boldsymbol{\theta}^{\prime}\right)$ as follows:
-$\alpha_{2}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{2}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi\left(\boldsymbol{\theta}^{\prime}\right)}{q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$ (6)
-where $q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)$ denotes the actual proposal density given by:
-$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right).$ (7)
-As argued in [6], compared with the standard MCMC, a single cycle of DA–MCMC can be slightly less efficient in terms of effective sample size per iteration.
-However, DA–MCMC can be more efficient in terms of the effective sample size per second because it does not have to perform the unnecessary computation of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)^{-1}$ and $\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)\right|$.
+To reduce computational costs, this study proposes to employ a DA–MCMC algorithm [5] which proceeds in two stages.
+We first approximate the target posterior distribution $\pi\left(\boldsymbol{\theta}\right)$ with a computationally cheaper surrogate $\pi^{*}\left(\boldsymbol{\theta}\right)$, which avoids evaluating the most expensive components—specifically, the inverse and determinant of $\boldsymbol{W}\left(\boldsymbol{\theta}\right)$.
+The approximate posterior is constructed by replacing $\boldsymbol{W}\left(\boldsymbol{\theta}\right)$ with the current value $\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)$, as follows:
+$\pi^{*}\left(\boldsymbol{\theta}\right)=\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)\right|^{-\frac{1}{2}}\exp\left\{-\frac{N}{2}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)\right\}p\left(\boldsymbol{\theta}\right).$
+In the first stage, a candidate $\boldsymbol{\theta}^{\prime}$ is proposed from a proposal distribution $q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)$.
+It is then preliminarily accepted with probability:
+$\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{1}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)}{q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi^{*}\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$
+If accepted, the candidate is promoted to the second stage, where it is evaluated under the full posterior $\pi\left(\boldsymbol{\theta}\right)$.
+Otherwise, it is rejected and the chain remains at the current state $\boldsymbol{\theta}^{\left(t+1\right)}=\boldsymbol{\theta}^{\left(t\right)}$.
+In the second stage, the final acceptance probability is computed as:
+$\alpha_{2}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{2}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi\left(\boldsymbol{\theta}^{\prime}\right)}{q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\},$
+where the effective second-stage proposal distribution is corrected as:
+$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right).$
+This structure improves computational efficiency by rejecting implausible proposals early, while still ensuring exact sampling from the full posterior distribution.
+As argued in [5], compared with the standard MCMC, a single cycle of DA–MCMC is slightly less efficient in terms of effective sample size per iteration.
+However, DA–MCMC can be more efficient in terms of the effective sample size per second because it does not require the unnecessary computation of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)^{-1}$ and $\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)\right|$.
 As shown in the next section, DA–MCMC is effective in our context.
-This study uses an adaptive random walk Metropolis–Hastings kernel, $q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\mathrm{N}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)},\;\varepsilon\boldsymbol{\Sigma}\right)$, where $\varepsilon>0$ denotes a step size, $\boldsymbol{\Sigma}$ denotes the $K$-dimensional symmetric positive definite matrix, and $\mathrm{N}\left(\boldsymbol{x}|\boldsymbol{a},\boldsymbol{B}\right)$ denotes a probability density function of a multivariate normal distribution with mean $\boldsymbol{a}$ and covariance matrix $\boldsymbol{B}$ evaluated at $\boldsymbol{x}$.
-The parameters of the proposal kernel are adaptively selected on the fly of an MCMC algorithm [8].
-At the $t$-th cycle, $\varepsilon$ is updated as $\log\varepsilon\leftarrow\log\varepsilon+t^{-\varsigma}\left(\bar{\alpha}^{\left(t\right)}-\alpha^{\dagger}\right)$, where $\bar{\alpha}^{\left(t\right)}$ denotes the average acceptance probability up to the $t$-th cycle, $\alpha^{\dagger}\in\left(0,1\right)$ denotes the target acceptance probability, and $\varsigma\in\left(0.5,1\right)$ denotes a tuning parameter.
+This study uses an adaptive random walk Metropolis–Hastings kernel specified as:
+$q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\mathrm{N}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)},\;\varepsilon\boldsymbol{\Sigma}\right),$
+where $\varepsilon>0$ denotes the step size, $\boldsymbol{\Sigma}$ denotes the $K$-dimensional symmetric positive definite matrix, and $\mathrm{N}\left(\boldsymbol{x}|\boldsymbol{a},\boldsymbol{B}\right)$ denotes the probability density function of a multivariate normal distribution with mean $\boldsymbol{a}$ and covariance matrix $\boldsymbol{B}$ evaluated at $\boldsymbol{x}$.
+The parameters of the proposal kernel are adaptively selected on the fly of the MCMC algorithm [7].
+At the $t$-th cycle, $\varepsilon$ is updated as $\log\varepsilon\leftarrow\log\varepsilon+t^{-\varsigma}\left(\bar{\alpha}^{\left(t\right)}-\alpha^{\dagger}\right)$, where $\bar{\alpha}^{\left(t\right)}$ denotes the average acceptance probability up to the $t$-th cycle, $\alpha^{\dagger}\in\left(0,1\right)$ denotes the target acceptance probability, and $\varsigma\in\left(0.5,1\right)$ denotes the tuning parameter.
 In this study, we fix $\varsigma=0.51$.
 $\boldsymbol{\Sigma}$ is updated based on the empirical covariance matrix of $\boldsymbol{\theta}^{\left(1\right)},\boldsymbol{\theta}^{\left(2\right)},...,\boldsymbol{\theta}^{\left(t\right)}$.
-The adaptation of $\varepsilon$ and $\boldsymbol{\Sigma}$ is performed only in the warmup phase and not thereafter.
+The adaptation of $\varepsilon$ and $\boldsymbol{\Sigma}$ is performed only during the warmup phase and not thereafter.
 3 Application
@@
 Acknowledgements
-This work was supported by JSPS KAKENHI Grant Number 20K22096.
+This work was supported by JSPS KAKENHI Grant Numbers 20K22096 and 25K21168.

### v3

@@
 Thus, the DA–MCMC algorithm offers a significant advancement in computational efficiency for quasi-Bayesian inference, making it a valuable tool for robust Bayesian analysis.
-…
 keywords:
@@
 1 Introduction
-In a standard Bayesian inference, a probabilistic model is used for the data-generating process and the corresponding likelihood function, $p\left(\mathcal{D}|\boldsymbol{\theta}\right)$, where $\mathcal{D}$ denotes a set of observations with a sample size of $N$, $\boldsymbol{\theta}\in\mathbb{R}^{K}$ denotes a $K$-dimensional vector of unknown parameters, and $p\left(\boldsymbol{\theta}\right)$ denotes a prior.
+In standard Bayesian inference, a probabilistic model is used for the data-generating process and the corresponding likelihood function, $p\left(\mathcal{D}|\boldsymbol{\theta}\right)$, where $\mathcal{D}$ denotes a set of observations with a sample size of $N$, $\boldsymbol{\theta}\in\mathbb{R}^{K}$ denotes a $K$-dimensional vector of unknown parameters, and $p\left(\boldsymbol{\theta}\right)$ denotes a prior.
 A posterior density is proportional to the product of the likelihood and prior, $p\left(\boldsymbol{\theta}|\mathcal{D}\right)\propto p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)$.
@@
 A well-known shortcoming of standard Bayesian analysis is that it needs to fully specify a probabilistic model.
-Alternative quasi-Bayesian approaches have been proposed to improve the robustness toward model misspecification, which have different constructions of quasi-likelihood that rely on loss function and moment conditions instead of a probabilistic model.
+Alternative quasi-Bayesian approaches have been proposed to improve the robustness toward model misspecification that rely on loss function and moment conditions instead of a probabilistic model.
 This study considers two types of quasi-Bayesian approaches.
@@
 Through applications to synthetic and real data problems, we demonstrate that the proposed algorithm is approximately two times more efficient than the standard MCMC in terms of the effective sample size per second.
-The remainder of this paper is organized as follows: Section 2 introduces DA–MCMC, Section 3 compares the proposed algorithm with existing algorithm via synthetic and real data examples, and Section 4 concludes the paper.
+The remainder of this paper is organized as follows: Section 2 introduces DA–MCMC, Section 3 compares the proposed algorithm with an existing algorithm via synthetic and real data examples, and Section 4 concludes the paper.
 2 Delayed acceptance MCMC
@@
 where the effective second-stage proposal distribution is corrected as:
-$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right).$
+$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)+\left(1-r\left(\boldsymbol{\theta}^{\left(t\right)}\right)\right)\delta_{\boldsymbol{\theta}^{\left(t\right)}}\left(\boldsymbol{\theta}^{\prime}\right),$
+with
+$r\left(\boldsymbol{\theta}^{\left(t\right)}\right)=\int\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)d\boldsymbol{\theta}^{\prime}.$
+Here $\delta_{\boldsymbol{\theta}^{\left(t\right)}}\left(\cdot\right)$ denotes the Dirac mass at $\boldsymbol{\theta}^{\left(t\right)}$.
 This structure improves computational efficiency by rejecting implausible proposals early, while still ensuring exact sampling from the full posterior distribution.
@@
 Table 1: Multivariate effective sample size per iteration
-Model $N$ $K$ multiESS/iter
+Model NN $K$ multiESS/iter
 MCMC DA–MCMC
@@
 Table 2: Multivariate effective sample size per second
-Model $N$ $K$ multiESS/s
+Model NN $K$ multiESS/s
 MCMC DA–MCMC
@@
 Table 3: Second-stage acceptance probability
-$N$ $K$ Percentile
+NN $K$ Percentile
 $P_{25}$ $P_{50}$ $P_{75}$